. What is the oxidation state of a. S in SO4- b. P in PO43 c. N in NH4+ d. S in H2S e. H in H2
Answer #1

The general oxidation state of

O is -2

H is +1

And the oxidation state of an element in its standard state is zero.

The sum of the oxidation states of all the elements in a compound or ion is equal to it's charge.

Let x be the oxidation state of required element.

Then

a. x + 4(-2)=-2

x= +6

So the oxidation state of S is +6

b. x +4(-2)=-3

x= +5

So the oxidation state of P is +5

c. x +4(+1)=+1

x=-3

So the oxidation state of N is -3

d. 2(+1)+x=0

x= -2

So the oxidation state of S is -2

e.H2 is the standard state so the oxidation state of H is 0